# thanks: Part of the Instituto Avanzado de Cosmología (IAC) Collaboration

### v1

@@ II.1 Previous Work and Unsolved Issues
Robles and Matos (2013a) studied the case of a real SF with mass $m$ and self-interaction $\lambda$, in a thermal bath with temperature $T$. Starting from the SF potential (1), it was shown that the characteristic polynomial for this system is given by
$\omega^{2}=k^{2}c^{2}+m^{2}c^{2}\left(1-\frac{T^{2}}{T_{c}^{2}}\right),$ (2)
@@
where $\omega$ corresponds to the angular frequency of the temporal part of the solution, $k$ to the wave-number of the radial part, $c$ is the speed of light and $T_{c}=2m/\sqrt{\lambda}$ is its critical phase-transition temperature.
@@
$m_{\Phi}^{2}=m^{2}\left(1-\frac{T^{2}}{T_{c}^{2}}\right).$ (3)
Consequently, galaxies of different sizes exhibit different effective masses $m_{\Phi}$. For example, if we assume that the mass of the scalar field is $m\sim 10^{-21}\text{ eV}/c^{2}$, which agrees with the constrictions of Lyman alpha observations (Linares Cedeño et al., 2021) and of the satellite galaxies of the Milky Way (Nadler et al., 2021), the effective SF mass is of the order of $m_{\Phi}\sim 10^{-24}\text{ eV}/c^{2}$. The halo of such a galaxy remains as a Bose-Einstein condensate, preserving its DM behavior. However, with the increase in temperature, some SFDM particles should occupy excited states and the effective mass obtained from different observations should be different, solving a current problem found with the Fuzzy DM model (see e.g. Ferreira, 2021; Bernal et al., 2018).
In (Robles and Matos, 2013a), the finite temperature density profile is obtained as
$\rho(r)=\sum_{j}\rho_{0}^{j}\frac{\sin^{2}(j\pi r/R)}{(j\pi r/R)^{2}},$ (4)
@@
The last multistate solution was shown capable to fit the rotation curves of low and high surface brightness galaxies and the wiggles observed in the rotation curves of some of them (Bernal et al., 2018), the strong gravitational lensing in SFDM halos (Robles and Matos, 2013c), reproduced successfully the evolution of a dwarf satellite galaxy embedded in the finite temperature SFDM halo in numerical simulations (Robles et al., 2015), and reproduced the dynamical mass distribution in clusters of galaxies (Bernal et al., 2017).
A disadvantage of this model is that different combinations of ground plus excited states are required to reproduce the observations, which could be explained due to the distinct formation processes of galaxies, groups and clusters of galaxies, resulting in different numbers of SF particles in diverse states.
II.2 Spontaneous Symmetry Breaking with a Complex Scalar Field
For the present article, we start with the following SF potential $V$ at early stages of the Universe, for a complex SF, $\Phi$:
$V=\frac{\lambda}{2}\left(\Phi\Phi^{*}-\frac{m^{2}}{\lambda}+\frac{1}{2}T^{2}\right)^{2},$ (5)
which is the same as potential (1) plus some gauge terms that depend on the SF mass $m$ and the temperature of the thermal bath $T$. In what follows, we will denote $\Phi\Phi^{*}:=n$ as the SF number density. From the first derivative of equation (5), we find that this potential has the following two extreme points:
$\begin{split}\Phi_{1}\Phi_{1}^{*}&=n_{1}=0;\\ \Phi_{2}\Phi_{2}^{*}&=n_{2}=\frac{m^{2}}{\lambda}-\frac{1}{2}T^{2}.\end{split}$ (6)
@@ II.2 Spontaneous Symmetry Breaking with a Complex Scalar Field
$V_{,\Phi\Phi^{*}}=\lambda\left(2n-\frac{m^{2}}{\lambda}+\frac{1}{2}T^{2}\right),$ (7)
which takes the value $V_{,\Phi\Phi^{*}}|_{n_{1}}=-m^{2}+\lambda T^{2}/2$ for the first critical point in (6), and $V_{,\Phi\Phi^{*}}|_{n_{2}}=m^{2}-\lambda T^{2}/2$ for the second one. If we define:
$T^{2}_{c}:=\frac{2m^{2}}{\lambda},$ (8)
the values of the second derivative (7) read $V_{,\Phi\Phi^{*}}=\pm(\lambda/2)\left(-T_{c}^{2}+T^{2}\right)$. This implies that the first extreme point is a minimum for $T^{2}>T_{c}^{2}$, and it is a maximum for $T^{2}<T_{c}^{2}$. The second extreme point does not exist for $T^{2}>T_{c}^{2}$ (since the corresponding $n_{2}<0$), and it is a double minimum for $T^{2}<T_{c}^{2}$ (see Figure 1). That means that $T_{c}$ is the critical temperature where the SSB takes place and the SF potential (5) changes from one minimum point to two minima and one maximum point. For the first critical value, $n_{1}=0$, the SF potential takes the value $V(n_{1})=(\lambda/2)\left(-\frac{m^{2}}{\lambda}+\frac{T^{2}}{2}\right)^{2}$, and for the second one, $n_{2}$, the SF potential is $V(n_{2})=0$. Observe that $n=0$ at the critical temperature $T_{c}$.
Figure 1: Qualitative behavior of the SF potential (5) for three values of the temperature $T$ with respect to the critical temperature $T_{c}$. The dotted blue line corresponds to $T>T_{c}$, where the system oscillates around the minimum of the potential. The dashed yellow line corresponds to $T=T_{c}$ where the system undergoes the SSB. The solid green line $T<T_{c}$ corresponds to the new potential with two minima and one maximum point
@@
$m_{\Phi}^{2}=\left\{\begin{matrix}-m^{2}+\frac{\lambda}{2}T^{2}=\frac{\lambda}{2}(T^{2}-T_{c}^{2})&\text{for}\,\,\,n_{1};\quad T^{2}>T_{c}^{2},\\ \\ m^{2}-\frac{\lambda}{2}T^{2}=\frac{\lambda}{2}(T_{c}^{2}-T^{2})&\text{for}\,\,\,n_{2};\quad T^{2}<T_{c}^{2},\end{matrix}\right.$ (9)
therefore $m_{\Phi}^{2}$ is always positive. As in the case of the real SF (equation (3)), notice from the last result that the effective mass of the complex SF depends on the temperature $T$ of the SF at the moment the halo collapsed, and the critical temperature $T_{c}$, which depends itself on the mass $m$ and the self-interaction parameter $\lambda$, intrinsic properties of the SF particle. This implies that the effective observed mass will depend on the particular properties of the observed system. Again, this may explain the tensions on the SF mass obtained from the fitting with diverse astrophysical and cosmological observations for the fuzzy DM model (see e.g. Ferreira, 2021; Bernal et al., 2018).
II.3 Evolution equations and Newtonian Limit
Now, we study the evolution of the SF with the potential (5) and finite temperature, in a FLRW Universe. The SF fulfills the Einstein-Klein-Gordon equations (EKG), $G_{\mu\nu}=\kappa^{2}T_{\mu\nu}$, and $\Box\Phi-\frac{dV}{d\Phi^{*}}=0$, where $G_{\mu\nu}$ is the Einstein tensor, $T_{\mu\nu}$ is the energy-momentum tensor of the SF and $\kappa^{2}:=8\pi G/c^{4}$. If we separate, as usual, the background part from a perturbed fluctuation, $\Phi=\Phi_{0}+\delta\Phi$, the fluctuation $\delta\Phi$ fulfills the following equations:
$\begin{split}&\nabla^{2}\delta\Phi-\ddot{\delta\Phi}-2H\dot{\delta\Phi}+V_{,\Phi_{0}\Phi_{0}^{*}}a^{2}\delta\Phi\\ &+V,_{\Phi^{*}_{0}\Phi^{*}_{0}}a^{2}\delta\Phi^{*}-2V,_{\Phi^{*}_{0}}a^{2}\phi+4\dot{\phi}\dot{\Phi}_{0}=0,\end{split}$ (10)
$\begin{split}&2\nabla^{2}\phi-6H(\dot{\phi}+H\phi)=\\ &\kappa^{2}\left[(\dot{\Phi}_{0}\dot{\delta\Phi}^{\ast}+\dot{\Phi}^{\ast}_{0}\dot{\delta\Phi})-2\phi\dot{\Phi}_{0}\dot{\Phi}^{\ast}_{0}+a^{2}\delta V\right],\end{split}$ (11)
@@ II.3 Evolution equations and Newtonian Limit
$\delta V:=V,_{\Phi_{0}}\delta\Phi+V,_{\Phi_{0}^{\ast}}\delta\Phi^{\ast}.$ (12)
Equation (10) is the KG equation for the perturbation $\delta\Phi$, and equation (11) is the Poisson equation.
@@
For the SF potential (5), equations (10) and (11) read
$\begin{split}\nabla^{2}\delta\Phi-\ddot{\delta\Phi}+\lambda\left(2n_{0}-\frac{m^{2}}{\lambda}+\frac{1}{2}T^{2}\right)a^{2}\delta\Phi&\\ +\lambda\Phi_{0}^{2}a^{2}\delta\Phi^{*}&=0,\end{split}$ (13)
$\begin{split}\nabla^{2}\phi=&\frac{\kappa^{2}}{2}\bigg{[}\left(\dot{\Phi}_{0}\dot{\delta\Phi}^{\ast}+\dot{\Phi}^{\ast}_{0}\dot{\delta\Phi}\right)-2\phi\dot{\Phi}_{0}\dot{\Phi}^{\ast}_{0}\\ &\left.+a^{2}\lambda\left(\Phi_{0}\Phi_{0}^{*}+\frac{m^{2}}{\lambda}-\frac{1}{2}T^{2}\right)\left(\Phi_{0}\delta\Phi^{*}+\Phi_{0}^{*}\delta\Phi\right)\right].\end{split}$ (14)
II.4 Hydrodynamic representation
To solve the last equations, we carry out a Madelung transformation:
$\Phi_{0}$ $=$ $\sqrt{n_{0}}\,e^{i\theta_{0}}=\sqrt{n_{0}}\,e^{i(S_{0}-\omega_{0}t)},$ (15)
$\delta\Phi$ $=$ $R\,e^{i(\theta_{0}+\delta S)}=\sqrt{n}\,e^{i(\theta_{0}-\omega_{0}t+\delta S)},$ (16)
where $\Phi$ is decomposed into a density number of the SF $n$ and a phase $S$. We will take perturbations such that if $\phi$ is of the order of $\epsilon$, then $\delta\Phi$, $\delta R$ and $\delta S$ are also of this order, while $n$ is a perturbation of order $\epsilon^{2}$. We will neglect perturbations of order $\epsilon^{2}$ and beyond. With these new variables, equations (13) and (14) transform into
$\nabla^{2}R-\ddot{R}$ (17)
$+\frac{\lambda}{2}\left(4n_{0}-T^{2}_{c}+T^{2}\right)a^{2}R+\dot{\theta}_{0}^{2}R-i(\dot{\theta}_{0}R^{2}\dot{)}=0,$
$\nabla^{2}\phi=\frac{\kappa^{2}}{2}\left[-\left(2\dot{\theta}_{0}^{2}n_{0}+\frac{\dot{n}_{0}^{2}}{2n_{0}}\right)\phi\right.$ (18)
$\left.+\frac{\dot{n}_{0}}{\sqrt{n_{0}}}\dot{R}+\left(\lambda(2n_{0}+T_{c}^{2}-T^{2})a^{2}+2\dot{\theta}_{0}^{2}\right)\sqrt{n_{0}}R\right].$
@@ II.4 Hydrodynamic representation
II.5 Exact solution
To solve the last field equations, we perform a separation of variables defining $R:=L_{k}Y_{k}^{j}T_{R}$, where $L_{k}=L_{k}(r)$, $T_{R}=T_{R}(\eta)$ and $Y_{k}^{j}=Y_{k}^{j}(\theta,\varphi)$ are the spherical harmonics polynomials, with $r$ the radial, $\eta$ the temporal, and $\theta$ and $\varphi$ the angular coordinates. The real part of equation (17) has the solution
$R=\sum_{k,j}\frac{J_{k+1/2}(l\,r)}{\sqrt{r}}Y^{j}_{k}T_{R},$ (19)
@@ II.5 Exact solution
$\ddot{T}_{R}+\left(\omega^{2}-2\lambda n_{0}a^{2}-\dot{\theta}_{0}^{2}\right)T_{R}=0,$ (20)
@@
$l^{2}=\omega^{2}-\frac{\lambda}{2}\left(T^{2}_{c}-T^{2}\right)a^{2}=\omega^{2}-m_{\Phi}^{2}a^{2},$ (21)
@@
where we used the definition of the SF mass (9). The last equation is crucial for what follows. This equation means that $l$ is a parameter that depends on the value of the effective mass $m_{\Phi}$ and the vibration energy $\omega$. We will see that $l$ determines the scale of the galaxy, that is, the scale $l$ depends strongly on the galaxy we are studying and this scale depends on the temperature and energy with which the galaxy ends after its formation collapse. In other words, the final state of the galaxy determines the temperature of the SFDM, and the scale of the galaxy is determined by the effective mass given by this temperature. By fitting the exact solutions to the rotation curves of galaxies, we can obtain a value for the scale $l$, but unfortunately we cannot determine the final temperature of the SFDM halo, we can only have the combination $\omega^{2}-m_{\Phi}^{2}a^{2}$.
@@
For equation (18), we can expand the Newtonian potential as $\phi=pT_{p}Y^{j}_{k}$, where again $p=p(r)$, $T_{p}=T_{p}(\eta)$ and $Y_{k}^{j}=Y_{k}^{j}(\theta,\varphi)$. Such equation transforms into
$\frac{1}{r^{2}}\left(r^{2}p_{,r}\right)_{,r}+\left(\Omega_{1}-\frac{k(k+1)}{r^{2}}\right)p=\Omega_{2}L_{k},$ (22)
where functions $\Omega_{i}=\Omega_{i}(\eta)$ are defined as
$\Omega_{1}=\frac{\kappa^{2}}{2}\left(2\dot{\theta}_{0}^{2}n_{0}+\frac{\dot{n}_{0}^{2}}{2n_{0}}\right),$ (23)
@@
$\Omega_{2}=\frac{\kappa^{2}}{2}\left[\frac{\dot{n}_{0}}{\sqrt{n_{0}}}\frac{\dot{T}_{R}}{T_{p}}+\left(\lambda(2n_{0}+T_{c}^{2}-T^{2})a^{2}+2\dot{\theta}_{0}^{2}\right)\sqrt{n_{0}}\frac{T_{R}}{T_{p}}\right].$ (24)
@@
Notice that the time dependence of the equations is due to the expansion of the Universe, which is too slow compared to the movements of the galaxies. Therefore, if we assume that such functions of the background are almost constant in time, as a good approximation we can set all these functions as constant. Thus, we can consider $\Omega_{1}$ and $\Omega_{2}$ as constants. In that case, equation (22) can be solved analytically. Furthermore, equation (20) for the time evolution of function $R$ is $T_{R}=T_{0}\exp(i(\omega^{2}-2\lambda n_{0}a^{2}-\dot{\theta}_{0}^{2})\eta)$, and the density number $n$ reduce to
$n=\sum_{k,j}\frac{J^{2}_{k+1/2}(l\,r)}{r}Y^{j}_{k}Y^{*j}_{k},$ (25)
@@
From equation (19), using the orthogonality of functions $Y^{j}_{k}$, we can define the complex function $R_{k}$ as
$R_{k}=\frac{J_{k+1/2}(l\,r)}{\sqrt{r}}\exp(i\omega_{k}\eta).$ (26)
@@
$R_{0}=\frac{J_{1/2}(lr)}{\sqrt{r}}\exp(i\omega_{0}\eta)=R_{00}\frac{\sin x}{x}\exp(i\omega_{0}\eta),$ (27)
$R_{1}=\frac{J_{3/2}(lr)}{\sqrt{r}}\exp(i\omega_{1}\eta)=R_{10}\left(\frac{\sin x}{x^{2}}-\frac{\cos x}{x}\right)\exp(i\omega_{1}\eta),$ (28)

### v3

@@ 2.1 Previous Work and Unsolved Issues
[43] studied the case of a real SF with mass $m$ and self-interaction $\lambda$, in a thermal bath with temperature $T$. Starting from the SF potential (2.1), it was shown that the characteristic polynomial for this system is given by
$\omega^{2}=k^{2}c^{2}+m^{2}c^{2}\left(1-\frac{T^{2}}{T_{c}^{2}}\right),$ (2.2)
@@
where $\omega$ corresponds to the angular frequency of the temporal part of the solution, $k$ to the wave number of the radial part, $c$ is the speed of light and $T_{c}=2m/\sqrt{\lambda}$ is its critical phase-transition temperature.
@@
$m_{\Phi}^{2}=m^{2}\left(1-\frac{T^{2}}{T_{c}^{2}}\right).$ (2.3)
Consequently, galaxies of different sizes exhibit different effective masses $m_{\Phi}$. For example, if we assume that the mass of the scalar field is $m\sim 10^{-21}\text{ eV}/c^{2}$, which agrees with the constrictions of Lyman alpha observations [51] and of the satellite galaxies of the Milky Way [52], the effective SF mass is of the order of $m_{\Phi}\sim 10^{-24}\text{ eV}/c^{2}$. The halo of such a galaxy remains as a Bose-Einstein condensate, preserving its DM behavior. However, with the increase in temperature, some SFDM particles should occupy excited states, and the effective mass obtained from different observations should be different, solving a current problem found with the Fuzzy DM model [see e.g. 53, 23].
In [43], the finite temperature density profile is obtained as
$\rho(r)=\sum_{j}\rho_{0}^{j}\frac{\sin^{2}(j\pi r/R)}{(j\pi r/R)^{2}},$ (2.4)
@@
The last multistate solution was shown capable of fitting the rotation curves of low and high surface brightness galaxies and the wiggles observed in the rotation curves of some of them [23], the strong gravitational lensing in SFDM halos [54], reproduced successfully the evolution of a dwarf satellite galaxy embedded in the finite temperature SFDM halo in numerical simulations [55] and reproduced the dynamical mass distribution in clusters of galaxies [56].
A disadvantage of this model is that different combinations of ground plus excited states are required to reproduce the observations, which could be explained due to the distinct formation processes of galaxies, groups, and clusters of galaxies, resulting in different numbers of SF particles in diverse states.
2.2 Spontaneous Symmetry Breaking with a Complex Scalar Field
For the present article, we start with the following SF potential $V$ at the early stages of the Universe, for a complex SF, $\Phi$:
$V=\frac{\lambda}{2}\left(\Phi\Phi^{*}-\frac{m^{2}}{\lambda}+\frac{1}{2}T^{2}\right)^{2},$ (2.5)
which is the same as potential (2.1) plus some gauge terms that depend on the SF mass $m$ and the temperature of the thermal bath $T$. We will denote $\Phi\Phi^{*}:=n$ as the SF number density in what follows. From the first derivative of equation (2.5), we find that this potential has the following two extreme points:
$\begin{split}\Phi_{1}\Phi_{1}^{*}&=n_{1}=0;\\ \Phi_{2}\Phi_{2}^{*}&=n_{2}=\frac{m^{2}}{\lambda}-\frac{1}{2}T^{2}.\end{split}$ (2.6)
@@ 2.2 Spontaneous Symmetry Breaking with a Complex Scalar Field
$V_{,\Phi\Phi^{*}}=\lambda\left(2n-\frac{m^{2}}{\lambda}+\frac{1}{2}T^{2}\right),$ (2.7)
which takes the value $V_{,\Phi\Phi^{*}}|_{n_{1}}=-m^{2}+\lambda T^{2}/2$ for the first critical point in (2.6), and $V_{,\Phi\Phi^{*}}|_{n_{2}}=m^{2}-\lambda T^{2}/2$ for the second one. If we define:
$T^{2}_{c}:=\frac{2m^{2}}{\lambda},$ (2.8)
the values of the second derivative (2.7) read $V_{,\Phi\Phi^{*}}=\pm(\lambda/2)\left(-T_{c}^{2}+T^{2}\right)$. This implies that the first extreme point is a minimum for $T^{2}>T_{c}^{2}$, and it is a maximum for $T^{2}<T_{c}^{2}$. The second extreme point does not exist for $T^{2}>T_{c}^{2}$ (since the corresponding $n_{2}<0$), and it is a double minimum for $T^{2}<T_{c}^{2}$ (see figure 1). That means that $T_{c}$ is the critical temperature where the SSB takes place and the SF potential (2.5) changes from one minimum point to two minima and one maximum point. For the first critical value, $n_{1}=0$, the SF potential takes the value $V(n_{1})=(\lambda/2)\left(-\frac{m^{2}}{\lambda}+\frac{T^{2}}{2}\right)^{2}$, and for the second one, $n_{2}$, the SF potential is $V(n_{2})=0$. Observe that $n=0$ at the critical temperature $T_{c}$.
Figure 1: Qualitative behavior of the SF potential (2.5) for three values of the temperature $T$ with respect to the critical temperature $T_{c}$. The dotted blue line corresponds to $T>T_{c}$, where the system oscillates around the minimum of the potential. The dashed yellow line corresponds to $T=T_{c}$ where the system undergoes the SSB. The solid green line $T<T_{c}$ corresponds to the new potential with two minima and one maximum point
@@
$m_{\Phi}^{2}=\left\{\begin{matrix}-m^{2}+\frac{\lambda}{2}T^{2}=\frac{\lambda}{2}(T^{2}-T_{c}^{2})&\text{for}\,\,\,n_{1};\quad T^{2}>T_{c}^{2},\\ \\ m^{2}-\frac{\lambda}{2}T^{2}=\frac{\lambda}{2}(T_{c}^{2}-T^{2})&\text{for}\,\,\,n_{2};\quad T^{2}<T_{c}^{2},\end{matrix}\right.$ (2.9)
therefore $m_{\Phi}^{2}$ is always positive. As in the case of the real SF (equation (2.3)), notice from the last result that the effective mass of the complex SF depends on the temperature $T$ of the SF at the moment the halo collapsed, and the critical temperature $T_{c}$, which depends itself on the mass $m$ and the self-interaction parameter $\lambda$, intrinsic properties of the SF particle. This implies that the effective observed mass will depend on the particular properties of the observed system. Again, this may explain the tensions on the SF mass obtained from the fitting with diverse astrophysical and cosmological observations for the fuzzy DM model [see bellow and e.g. 53, 23].
2.3 Evolution Equations and Newtonian Limit
Now, we study the evolution of the SF with the potential (2.5) and finite temperature, in a FLRW Universe. The SF fulfills the Einstein-Klein-Gordon equations (EKG), $G_{\mu\nu}=\kappa^{2}T_{\mu\nu}$, and $\Box\Phi-\frac{dV}{d\Phi^{*}}=0$, where $G_{\mu\nu}$ is the Einstein tensor, $T_{\mu\nu}$ is the energy-momentum tensor of the SF and $\kappa^{2}:=8\pi G/c^{4}$. If we separate, as usual, the background part from a perturbed fluctuation, $\Phi=\Phi_{0}+\delta\Phi$, the fluctuation $\delta\Phi$ fulfills the following equations:
$\nabla^{2}\delta\Phi-\ddot{\delta\Phi}-2H\dot{\delta\Phi}+V_{,\Phi_{0}\Phi_{0}^{*}}a^{2}\delta\Phi+V,_{\Phi^{*}_{0}\Phi^{*}_{0}}a^{2}\delta\Phi^{*}-2V,_{\Phi^{*}_{0}}a^{2}\phi+4\dot{\phi}\dot{\Phi}_{0}=0,$ (2.10)
$2\nabla^{2}\phi-6H(\dot{\phi}+H\phi)=\kappa^{2}\left[(\dot{\Phi}_{0}\dot{\delta\Phi}^{\ast}+\dot{\Phi}^{\ast}_{0}\dot{\delta\Phi})-2\phi\dot{\Phi}_{0}\dot{\Phi}^{\ast}_{0}+a^{2}\delta V\right],$ (2.11)
@@ 2.3 Evolution Equations and Newtonian Limit
$\delta V:=V,_{\Phi_{0}}\delta\Phi+V,_{\Phi_{0}^{\ast}}\delta\Phi^{\ast}.$ (2.12)
Equation (2.10) is the KG equation for the perturbation $\delta\Phi$, and equation (2.11) is the Poisson equation.
@@
For the SF potential (2.5), equations (2.10) and (2.11) read
$\nabla^{2}\delta\Phi-\ddot{\delta\Phi}+\lambda\left(2n_{0}-\frac{m^{2}}{\lambda}+\frac{1}{2}T^{2}\right)a^{2}\delta\Phi+\lambda\Phi_{0}^{2}a^{2}\delta\Phi^{*}=0,$ (2.13)
$\nabla^{2}\phi=\frac{\kappa^{2}}{2}\left[\left(\dot{\Phi}_{0}\dot{\delta\Phi}^{\ast}+\dot{\Phi}^{\ast}_{0}\dot{\delta\Phi}\right)-2\phi\dot{\Phi}_{0}\dot{\Phi}^{\ast}_{0}+a^{2}\lambda\left(\Phi_{0}\Phi_{0}^{*}+\frac{m^{2}}{\lambda}-\frac{1}{2}T^{2}\right)\left(\Phi_{0}\delta\Phi^{*}+\Phi_{0}^{*}\delta\Phi\right)\right].$ (2.14)
2.4 Hydrodynamic Representation
To solve the last equations, we carry out a Madelung transformation [57]:
$\Phi_{0}$ $=$ $\sqrt{n_{0}}\,e^{i\theta_{0}}=\sqrt{n_{0}}\,e^{i(S_{0}-\omega_{0}t)},$ (2.15)
$\delta\Phi$ $=$ $R\,e^{i(\theta_{0}+\delta S)}=\sqrt{n}\,e^{i(\theta_{0}-\omega_{0}t+\delta S)},$ (2.16)
where $\Phi$ is decomposed into a density number of the SF $n$ and a phase $S$. We will take perturbations such that if $\phi$ is of the order of $\epsilon$, then $\delta\Phi$, $\delta R$, and $\delta S$ are also of this order, while $n$ is a perturbation of order $\epsilon^{2}$. We will neglect perturbations of order $\epsilon^{2}$ and beyond. With these new variables, equations (2.13) and (2.14) transform into
$\nabla^{2}R-\ddot{R}+\frac{\lambda}{2}\left(4n_{0}-T^{2}_{c}+T^{2}\right)a^{2}R+\dot{\theta}_{0}^{2}R-i(\dot{\theta}_{0}R^{2}\dot{)}=0,$ (2.17)
$\nabla^{2}\phi=\frac{\kappa^{2}}{2}\left[-\left(2\dot{\theta}_{0}^{2}n_{0}+\frac{\dot{n}_{0}^{2}}{2n_{0}}\right)\phi+\frac{\dot{n}_{0}}{\sqrt{n_{0}}}\dot{R}+\left(\lambda(2n_{0}+T_{c}^{2}-T^{2})a^{2}+2\dot{\theta}_{0}^{2}\right)\sqrt{n_{0}}R\right].$ (2.18)
@@ 2.4 Hydrodynamic Representation
2.5 Exact Solution
To solve the last field equations, we perform a separation of variables defining $R:=L_{k}Y_{k}^{j}T_{R}$, where $L_{k}=L_{k}(r)$, $T_{R}=T_{R}(\eta)$ and $Y_{k}^{j}=Y_{k}^{j}(\theta,\varphi)$ are the spherical harmonics polynomials, with $r$ the radial, $\eta$ the temporal, and $\theta$ and $\varphi$ the angular coordinates. The real part of equation (2.17) has the solution
$R=\sum_{k,j}\frac{J_{k+1/2}(l\,r)}{\sqrt{r}}Y^{j}_{k}T_{R},$ (2.19)
@@ 2.5 Exact Solution
$\ddot{T}_{R}+\left(\omega^{2}-2\lambda n_{0}a^{2}-\dot{\theta}_{0}^{2}\right)T_{R}=0,$ (2.20)
@@
$l^{2}=\omega^{2}-\frac{\lambda}{2}\left(T^{2}_{c}-T^{2}\right)a^{2}=\omega^{2}-m_{\Phi}^{2}a^{2},$ (2.21)
@@
where we used the definition of the SF mass (2.9). The last equation is crucial for what follows. This equation means that $l$ is a parameter that depends on the value of the effective mass $m_{\Phi}$ and the vibration energy $\omega$. We will see that $l$ determines the scale of the galaxy, i.e. the scale $l$ depends strongly on the galaxy we are studying and this scale depends on the temperature and energy with which the galaxy ends after its formation collapse. In other words, the final state of the galaxy determines the temperature of the SFDM, and the scale of the galaxy is determined by the effective mass given by this temperature. By fitting the exact solutions to the rotation curves of galaxies, we can obtain a value for the scale $l$, but unfortunately, we cannot determine the final temperature of the SFDM halo, we can only have the combination $\omega^{2}-m_{\Phi}^{2}a^{2}$. Thus, we can say that equation (2.21) can solve the mass tension between the mass that fits Lyman-$\alpha$ observations and the mass needed to fit rotation curves in galaxies. Let us explain this point a bit further. It has been argued that the mass of the SF $m_{\Phi}$ appears to be different to fit observations of large galaxies, small galaxies, and observations of Lyman-$\alpha$ [31, 58, 59, 60, 61, 62, 63, 64]. It is clear from equation (2.9) that after the turn-around of the galactic halo occurs, the SFDM changes its temperature because it collapses again. Depending on the fluctuation size, environmental conditions, etc., a final SFDM halo temperature will be reached, and hence an effective frequency $\omega$ and the effective SF mass $m_{\Phi}$. There are two possibilities here. The first is that the self-interactions parameter is of order $\lambda\sim 1$, and then if $T\sim T_{c}$, the SFDM mass $m_{\Phi}$ could be the one observed using Lyman-$\alpha$ observations, $m_{\Phi}\sim 10^{-21}$eV, and the effective mass of a medium-sized galaxy given by the equation (2.9) could be $m_{\Phi}\sim 10^{-22}$eV. The second possibility is that the self-interaction parameter $\lambda\ll 1$, as claimed in [50]. In that case, $T^{2}/T_{c}^{2}\sim 0$ and it is sufficient that $m_{\Phi}\sim\omega$ for the scale $l$ to fit well to any galaxy. In other words, equation (2.21) can relieve the mass tension of the SFDM model, and the scale $l$ of a galaxy is determined not only by the mass $m_{\Phi}$ but also by the parameter $\omega$.
@@
For equation (2.18), we can expand the Newtonian potential as $\phi=pT_{p}Y^{j}_{k}$, where again $p=p(r)$, $T_{p}=T_{p}(\eta)$ and $Y_{k}^{j}=Y_{k}^{j}(\theta,\varphi)$. Such an equation transforms into
$\frac{1}{r^{2}}\left(r^{2}p_{,r}\right)_{,r}+\left(\Omega_{1}^{2}-\frac{k(k+1)}{r^{2}}\right)p=\Omega_{2}^{2}L_{k},$ (2.22)
whose solution is
$p=\sum_{k,j}\left(\frac{J_{k+1/2}(\Omega_{1}\,r)}{\sqrt{r}}+\Omega_{2}^{2}p_{0}\frac{J_{k+1/2}(l\,r)}{\sqrt{r}(\Omega_{1}^{2}-l^{2})}\right)$ (2.23)
where $p_{0}$ is an integration constant. It can be seen that solution (2.23) is well behaved even at $\Omega_{1}=l$. Functions $\Omega_{i}=\Omega_{i}(\eta)$ are defined as
$\Omega_{1}^{2}=\frac{\kappa^{2}}{2}\left(2\dot{\theta}_{0}^{2}n_{0}+\frac{\dot{n}_{0}^{2}}{2n_{0}}\right)=\kappa^{2}\dot{\Phi}_{0}\dot{\Phi}_{0}^{*},$ (2.24)
@@
$\Omega_{2}^{2}=\frac{\kappa^{2}}{2}\left[\frac{\dot{n}_{0}}{\sqrt{n_{0}}}\frac{\dot{T}_{R}}{T_{p}}+\left(\lambda(2n_{0}+T_{c}^{2}-T^{2})a^{2}+2\dot{\theta}_{0}^{2}\right)\sqrt{n_{0}}\frac{T_{R}}{T_{p}}\right].$ (2.25)
One interesting solution is when
$T_{p}=\frac{\kappa^{2}}{2}\left[\frac{\dot{n}_{0}}{\sqrt{n_{0}}}{\dot{T}_{R}}+\left(\lambda(2n_{0}+T_{c}^{2}-T^{2})a^{2}+2\dot{\theta}_{0}^{2}\right)\sqrt{n_{0}}{T_{R}}\right],$ (2.26)
@@
in that case $\Omega_{2}=1$. Notice that the time dependence of the equations is due to the expansion of the Universe, which is too slow compared to the movements of the galaxies. Therefore, if we assume that such functions of the background are almost constant in time, as a good approximation we can set all these functions as constant. Thus, we can consider $\Omega_{1}$ and $\Omega_{2}$ as constants. In that case, equation (2.22) can be solved analytically. Furthermore, equation (2.20) for the time evolution of function $R$ is $T_{R}=T_{0}\exp(i(\omega^{2}-2\lambda n_{0}a^{2}-\dot{\theta}_{0}^{2})\eta)$, and the density number $n$ reduce to
$n=\sum_{k,j}\frac{J^{2}_{k+1/2}(l\,r)}{r}Y^{j}_{k}Y^{*j}_{k},$ (2.27)
@@
From equation (2.19), using the orthogonality of functions $Y^{j}_{k}$, we can define the complex function $R_{k}$ as
$R_{k}=\frac{J_{k+1/2}(l\,r)}{\sqrt{r}}\exp(i\omega_{k}\eta).$ (2.28)
@@
$R_{0}=\frac{J_{1/2}(lr)}{\sqrt{r}}\exp(i\omega_{0}\eta)=R_{00}\frac{\sin x}{x}\exp(i\omega_{0}\eta),$ (2.29)
$R_{1}=\frac{J_{3/2}(lr)}{\sqrt{r}}\exp(i\omega_{1}\eta)=R_{10}\left(\frac{\sin x}{x^{2}}-\frac{\cos x}{x}\right)\exp(i\omega_{1}\eta),$ (2.30)